Let f be a differentiable function with f(0)=-4 and f(10)=11. which of the following must be true for some c in the interval (0,10)?


A) f'(c)=0, since the Extreme Value Theorem applies

B) f'(c)=11+(-4)/10-0 since the Mean Value Theorem applies

C) f'(c)=11-(-4)/10-0, since the Mean Value Theorem applies

D) f'(c)=1.5, since the Intermediate Value Theorem applies


1
Expert's answer
2021-01-19T18:26:32-0500

The mean value theorem states that if "f" is a continuous function on the closed interval "[a,b]" and differentiable on the open interval "(a,b)" , then there exists a point "c" in "(a,b)" such that

"f'(c)=\frac{f(b)-f(a)}{b-a}"


We have:

"\frac{f(b)-f(a)}{b-a}=\frac{11-(-4)}{10-0}=f'(c)" , "c\isin (0,10)"


Answer: C) f'(c)=11-(-4)/10-0, since the Mean Value Theorem applies